40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Fri Nov 4 23:34:28 2016, awade, DailyProgress, Electronics Equipment, Trouble shooting in the north path 100_0258.JPG100_0261.JPG
    Reply  Tue Nov 8 14:01:11 2016, awade, DailyProgress, Electronics Equipment, Trouble shooting in the north path 2016-11-08_11.35.34.jpg
       Reply  Wed Nov 16 12:46:59 2016, awade, DailyProgress, Electronics Equipment, Trouble shooting in the north path and Murphy's law 
          Reply  Wed Dec 7 18:59:46 2016, awade, DailyProgress, Electronics Equipment, FSS monitor daughter board damage and Murphy's law D040424-B.pdf
             Reply  Wed Dec 7 21:14:07 2016, Koji, DailyProgress, Electronics Equipment, FSS monitor daughter board damage and Murphy's law 
                Reply  Thu Dec 8 17:37:32 2016, awade, DailyProgress, Electronics Equipment, FSS monitor daughter board fixed 
       Reply  Thu Jan 5 15:50:55 2017, awade, DailyProgress, Electronics Equipment, Trouble shooting in the north path 2017-01-05_15.17.55.jpg
Message ID: 1784     Entry time: Wed Dec 7 18:59:46 2016     In reply to: 1770     Reply to this: 1786
Author: awade 
Type: DailyProgress 
Category: Electronics Equipment 
Subject: FSS monitor daughter board damage and Murphy's law 

Following up on the possible damage to the common path daughter board, it appears that there is an issue. This board, DCC-D040424, buffers and then provides some amplification if one chooses to solder some resisters and capacitors in. Right now it should be simply buffering but it is railing to a fixed offset in the mixer circuit.  

The north (RCAV) field box was switched out for the south path (ACAV) box (that I know is working correctly) and I found it worked as expected.  I then switch the south box's daughter board into the North (RCAV) field box and it operated as normal. The issue is therefore in the daughter board only.

The mixer op amp circuit uses U1 and U3, neither of these chips have visible damage. U4 was the one with obvious heat damage from the previous short. As it had the most obvious damage, I desoldered U4 and changed it out for another LT1125CSW. This isn't easy as its surface mounted and looks like some kind of machine done job.  I think I did a little damage to the pads, but there is enough left to give a good enough electrical connection.  I need to check with Koji, or someone else with wisdom, what the best soldering iron temperature is for doing this kind of stuff. 

Plugging the daughter board back into the north field box, it appears that the channels associated with U4 are working fine, I will need to maybe replace the other LT1125 and OP27s tomorrow. We kind of need the monitor channels to diagnose issues and also for the PID laser slow controls so this needs to be fixed.

Do we have spares of these boards are they hard to reorder in lots of one?

 

Attachment 1: D040424-B.pdf  20 kB  | Hide | Hide all
D040424-B.pdf
ELOG V3.1.3-